10 March 2020 UT SOUL-C Night 2

Closed at sunset for clouds.

~02:00 Checked pupil and FS alignment. Both were slightly misaligned and an adjustment was made to align them.

02:30 Shane is putting the ARGOS swing arm in.

Closed dome tests of ESM mode

Shane and Alfio tested SW changes to remove two unnecessary steps (check flux and center pupils) in closing the loop with SOUL in ESM mode. Success.

05:55 Opening --- we'll start with GOpt test on a bright star ~R=8. Seeing on guider ~1.2"

06:20 Closing.

07:52 Starting data collection for NCPA measurements.
  • subwin 256 x 256 (SW 1 839 556 256 256) (DIT=0.3133s)
  • H, N30_FS.
09:25 At 330 deg, source not in subwin, too high.
  • new subwin (SW 1 910 765 256 256)
  • check position and focus on new image, 0073. Looks OK. Peak counts 3400.
~10:15 Using the Z5 and Z6 determined last night to maximize Strehl. A few checks at different rot angles were made earlier on the symmetric modes 4 and 11 and these look like they do not change with rot angle as expected. After that, Doug launched a script to step through Z5, Z6, Z7, Z8, Z9, Z10 at rotator angles 360,330,330...0

11:15 Safe-skipping. Loops opened. Rotangle 120.

11:16 Restarted.

12:12 Shane noticed that the loop is not looking good. We are seeing vibrations and the source image is very elongated. At rotangle 0.

12:24 Script stopped. Taking a few more points at different rot angles close to 0. As soon as moved the rotator, the vibrations stopped, but the gains went to zero. Now that they gains are back up, it looks fine. Rotangle 60.

12:31 Did a few measurements at rotangle 60, then 30. These were OK, but once again at 0, the vibrations started. Paused loop and still we see the vibrations. Doug moved to rotangle 2, still see vibrations. At rotangle 10, vibrations seen. Shane turned off high order gains & tip tilt down (what the script does when moving the rotator) and it settled. After re-optimizing, things have settled - gains are the same, but gopt is off.

12:42 Starting script at rotangle 0 to do all 5-10 zernikes, now that vibrations are gone. Script will do rotangles 0, 30 and 60.

12:25 Done.
  • Shane trying to park rotator at 341, but it did not move. (It was moving as commanded by the script. David is looking at it).
  • I stopped/started Instrument Manager on LUCI1 - custom masks do not appear.

-- OlgaKuhn - 10 Mar 2020
Topic revision: r4 - 10 Mar 2020, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback